OpenStudy (anonymous):

Will someone please help me.. I don't know what i did wrong with a couple of problems please help

OpenStudy (anonymous):

I'll help with one.

OpenStudy (anonymous):

okay it says to find the volume of the solid

OpenStudy (anonymous):

Do you know the formula for the volume of a cone?

OpenStudy (anonymous):

not of the top of my head

OpenStudy (anonymous):

sorry, my computer just froze up it is (1/3)*Pi*r^2*h

OpenStudy (anonymous):

are 10 and 18 the heights, or are they the length of the lines they are next to?

OpenStudy (anonymous):

yes so \[1/3\times \pi \times 5^{2}\times10\]

OpenStudy (anonymous):

only if 10 is the height

OpenStudy (anonymous):

\[1/3\times \pi \times 5^{2}\times18\]

OpenStudy (anonymous):

10 is the height of the top cone and 18 is the height of the bottom cone

OpenStudy (anonymous):

Actually, I don't think so. If you do it that way you get 733.04, which got counted wrong on your test.

OpenStudy (anonymous):

I doubt you would get it wrong for being off by .04

OpenStudy (anonymous):

So, lets assume that 18 and 10 are the lengths of the lines they are next to, and not the heights.

OpenStudy (anonymous):

You following so far?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Awesome. We need the heights to use our formula, so now we just need to figure out how to get them. Thinking in 3D is rather tricky, so I like to make this into something I can do in 2D.

OpenStudy (anonymous):

Notice how we have the radius and the hypotenuse of a right triangle?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

Well, we can use the good old Pythagorean formula to figure out the missing side.

OpenStudy (anonymous):

Sqrt[18^2-5^2] and Sqrt[10^2-5^2] are our two heights

OpenStudy (anonymous):

(1/3)Pi*( Sqrt[18^2-5^2] )^2 + (1/3)Pi*( Sqrt[10^2-5^2] )^2

OpenStudy (anonymous):

^ That should give you the correct answer, which comes out to be...

OpenStudy (anonymous):

idk i dont have a calculator

OpenStudy (anonymous):

oops, I did that formula wrong :/

OpenStudy (anonymous):

(1/3)Pi*( Sqrt[18^2-5^2] )^2 + (1/3)Pi*( Sqrt[10^2-5^2] )^2 isn't right, but you get the idea

OpenStudy (anonymous):

I left out the radius

OpenStudy (anonymous):

You get it though, right?

OpenStudy (anonymous):

kinda

OpenStudy (anonymous):

(1/3)*Pi*r^2*h

OpenStudy (anonymous):

you did it right, if 10 and 18 had been the height. Unfortunatly, they were the side length (confusing problem)

OpenStudy (anonymous):

you just use the Pythagorean formula to get the heights, and do the same thing with them that you did with 10 and 18 the first time around.

OpenStudy (anonymous):

Well, I'm sleepy, so I'm off to bed. I have to wake up and do calculus D: